Circular No. 5699

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993C IN NGC 2954 C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of an

apparent supernova of mag ∼ 18 in the elliptical galaxy NGC 2954, found on a Kodak Technical Pan film taken Jan. 27.04 UT and confirmed on a similar film taken Jan. 28.03 by D. Albanese. SN 1993C is located at α = 9^{h}37^{m}39^{s}.70, δ = +15^{o}09′49′′.8 (equinox 1950.0), which is 4′′.8 west and 49′′.7 north of the galaxy’s center. Nearby stars have the following positional end figures: 40^{s}.83, 48′′.6 (V = 17.6); 38^{s}.71, 10′30′′.0 (V ∼ 14).

S. Benetti and E. Cappellaro, Asiago Astrophysical Observatory, confirm the presence of the new object, citing the offsets as 3′′ west, 50′′ north; with the 1.82-m telescope (+ CCD camera) at Cima Ekar, they obtained the following magnitudes, which they note are consistent with a type-Ia supernova ∼ 1 month after maximum light: Jan. 27.92 UT, B = 18.7, V = 17.7, R = 17.4.

SUPERNOVAE 1993A and 1993B IN ANONYMOUS GALAXIES M. M. Phillips, Cerro Tololo Interamerican Observatory, reports that

a spectrogram of SN 1993A (cf. IAUC 5693) obtained with the CTIO 4.0m telescope on Jan. 28.22 UT reveals strong, broad Hα emission with the characteristic P-Cyg profile of a type-II plateau event. Weaker, broad absorption features due to Hβ and Fe II are also clearly visible. A spectrum obtained of the host galaxy yields a redshift of 0.028.

A spectrogram of SN 1993B obtained by Phillips on Jan. 28.28 UT shows that this object is a type-Ia supernova ∼ 10 days past maximum. The host galaxy redshift was measured to be 0.069.

COMET MUELLER (1993a) Total visual magnitude estimates: Jan. 16.31 UT, 13.4 (A. Hale, Las

Cruces, NM, 0.41-m reflector); 19.01, 13.6 (H. Mikuz, Ljubljana, Slovenia, 0.20-m Baker-Schmidt camera + CCD + V filter; 5′ fan-shaped tail in p.a. 160^{o}); 21.60, 13.3 (A. Nakamura, Kuma, Japan, 0.60-m reflector); 25.91, 14.0 (Mikuz, 0.19-m f/4 flat field camera + CCD + V filter).

PERIODIC COMET VÄISÄLÄ 1 (1992u) Total visual magnitude estimate by Nakamura: Jan. 21.65 UT, 15.0.

1993 January 28 Daniel W. E. Green